• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 22:22:59 on June 05, 2010

    Entry number 324755

    keyword=PREX experiment status Saturday June 5th 2010 - No beam for now

    We got beam back after the CHL crash at the beginning of owl shift. Slug 28 was complete with about 1.7 C the total accumulated charge so far being around 31.7 C. The beam was tuned through the chicane and signal over backgroud ratio seems good enough to make a measurement. Unfortunately the CHL tripped again, and cryo as been working on resuming it. So far we do not expect beam to come during swing shift, so this swing shift and owl shift are on stand by, only the target operator needs to be there to monitor the target and call the 3rd and Shift Leader when the beam will start to be resumed. So if you are on Sunday Owl shift please send me your contact numbers so we can reach you when the beam comes back.

    Regards,

    Alexandre Camsonne Current PREX Hall A run coordinator